Please help me solve this equation:
what is one third of 3 to the ninth power?

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

what is one third of 3 to the ninth power?
---
Ans: (1/3)3^9 = 3*8 = 6561
===========================
Cheers,
Stan H.
=============

Answer by mananth(16946)